There are 30 people total (Men,women,children) There are 5 more men than women. There are x amount of women. What represents the total number of children.
A. 25-x
B. 35-2x
c. 25-2x
d. other

Answer by richwmiller(17219)

p=30
30=m+x+c
m=5+x
30=5+x+x+c
25-2x=c
that would be c.)